\begin{thebibliography}{99}

\bibitem{Whee} R. Ruffini and J. A. Wheeler, Phys. Today {\bf 24}, 30
(1971).

\bibitem{Car} B. Carter, in {\it Black Holes}, Proceedings of 1972 Session of Ecole d'ete de Physique Theorique,
edited by C. De Witt and B. S. De Witt (Gordon and Breach, New York,
1973).

\bibitem{Bek1} A. E. Mayo and J. D. Bekenstein, Phys. Rev. D {\bf 54}, 5059 (1996).

\bibitem{Chas} J. E. Chase, Commun. Math. Phys. {\bf 19}, 276 (1970); J. D. Bekenstein,
Phys. Rev. Lett. {\bf 28}, 452 (1972); C. Teitelboim, Lett. Nuovo
Cimento {\bf 3}, 326 (1972); I. Pena and D. Sudarsky, Class. Quant.
Grav. {\bf 14}, 3131 (1997).

\bibitem{BekVec} J. D. Bekenstein, Phys. Rev. D {\bf 5}, 1239 (1972); {\bf 5}, 2403 (1972);
M. Heusler, J. Math. Phys. {\bf 33}, 3497 (1992); D. Sudarsky,
Class. Quant. Grav. {\bf 12}, 579 (1995).

\bibitem{Hart} J. Hartle, Phys. Rev. D {\bf 3}, 2938 (1971); C. Teitelboim, Lett.
Nuovo Cimento {\bf 3}, 397 (1972).

\bibitem{Nun} D. N\'u\~nez, H. Quevedo, and D. Sudarsky, Phys. Rev. Lett. {\bf 76}, 571 (1996).

\bibitem{Hod11} S. Hod, Phys. Rev. D {\bf 84}, 124030 (2011) [arXiv:1112.3286].

\bibitem{Hodrc} S. Hod, Phys. Rev. D {\bf 86}, 104026 (2012) [arXiv:1211.3202];
S. Hod, The Euro. Phys. Journal C {\bf 73}, 2378 (2013)
[arXiv:1311.5298]; S. Hod, Phys. Rev. D {\bf 90}, 024051 (2014)
[arXiv:1406.1179]; S. Hod, Phys. Lett. B {\bf 739}, 196 (2014)
[arXiv:1411.2609]; S. Hod, Class. and Quant. Grav. {\bf 32}, 134002
(2015); S. Hod, Class. and Quant. Grav. {\bf 33}, 114001 (2016); S.
Hod, Phys. Lett. B {\bf 758}, 181 (2016); S. Hod and O. Hod, Phys.
Rev. D {\bf 81}, 061502 Rapid communication (2010)
[arXiv:0910.0734]; S. Hod, Phys. Lett. B {\bf 708}, 320 (2012)
[arXiv:1205.1872].

\bibitem{HerR} C. A. R. Herdeiro and E. Radu, Phys. Rev. Lett. {\bf 112}, 221101
(2014); C. A. R. Herdeiro and E. Radu, Phys. Rev. D {\bf 89}, 124018
(2014); C. A. R. Herdeiro and E. Radu, Int. J. Mod. Phys. D {\bf
23}, 1442014 (2014); C. L. Benone, L. C. B. Crispino, C. Herdeiro,
and E. Radu, Phys. Rev. D {\bf 90}, 104024 (2014); C. Herdeiro, E.
Radu, and H. R\'unarsson, Phys. Lett. B {\bf 739}, 302 (2014); C.
Herdeiro and E. Radu, Class. Quantum Grav. {\bf 32} 144001 (2015);
C. A. R. Herdeiro and E. Radu, Int. J. Mod. Phys. D {\bf 24},
1542014 (2015); C. A. R. Herdeiro and E. Radu, Int. J. Mod. Phys. D
{\bf 24}, 1544022 (2015); J. C. Degollado and C. A. R. Herdeiro,
Gen. Rel. Grav. {\bf 45}, 2483 (2013); P. V. P. Cunha, C. A. R.
Herdeiro, E. Radu, and H. F. R\'unarsson, Phys. Rev. Lett. {\bf
115}, 211102 (2015); C. A. R. Herdeiro, E. Radu, and H. F.
R\'unarsson, Phys. Rev. D {\bf 92}, 084059 (2015); Y. Brihaye, C.
Herdeiro, and E. Radu, arXiv:1605.08901 .

\bibitem{Notetim} That is, the timescale it takes for the black hole to swallow
or to scatter away to infinity the external matter fields.

\bibitem{QNMs} H. P. Nollert, Class. Quant. Grav. {\bf 16}, R159 (1999);
E. Berti, V. Cardoso and A. O. Starinets, Class. Quant. Grav. {\bf
26}, 163001 (2009); C. V. Vishveshwara, Nature {\bf 227}, 936
(1970); E. W. Leaver, Proc. R. Soc. A {\bf 402}, 285 (1985); B.
Mashhoon, Phys. Rev. D {\bf 31}, 290 (1985); H. P. Nollert, Phys.
Rev. D {\bf 47}, 5253 (1993); S. Hod, Phys. Rev. Lett. {\bf 81},
4293 (1998) [arXiv:gr-qc/9812002]; G. T. Horowitz and V. E. Hubeny,
Phys. Rev. D {\bf 62}, 024027 (2000); K. Glampedakis and N.
Andersson, Phys. Rev. D {\bf 64}, 104021 (2001); S. Hod, Phys. Rev.
D {\bf 67}, 081501 (2003) [arXiv:gr-qc/0301122]; S. Hod and U.
Keshet, Class. Quant. Grav. {\bf 22}, L71 (2005)
[arXiv:gr-qc/0505112]; S. Hod and U. Keshet, Phys.Rev. D {\bf 73},
024003 (2006) [arXiv:hep-th/0506214]; U. Keshet and S. Hod, Phys.
Rev. D {\bf 76}, R061501 (2007) [arXiv:0705.1179]; A. Gruzinov,
arXiv:gr-qc/0705.1725; V. Cardoso, A. S. Miranda, E. Berti, H.
Witek, and V. T. Zanchin, Phys. Rev. D {\bf 79}, 064016 (2009); R.
A. Konoplya and A. Zhidenko, Rev. Mod. Phys. {\bf 83}, 793 (2011);
Y. D\'ecanini, A. Folacci, and B. Raffaelli, Phys. Rev. D {\bf 84},
084035 (2011); H. Yang, A. Zimmerman, A. Zenginoglu, F. Zhang, E.
Berti, and Yanbei Chen, Phys. Rev. D {\bf 88}, 044047 (2013); S.
Hod, Phys. Rev. D {\bf 88}, 084018 (2013) [arXiv:1311.3007].

\bibitem{Notetails} It is worth noting that, in asymptotically flat black-hole spacetimes,
these exponentially damped quasinormal oscillations, which
characterize the composed black-hole-field system, are followed by
decaying power-law wave tails \cite{Tails} which reflect the
backscattering of the fields by the effective curvature potential of
the black-hole spacetime at large distances from the central black
hole.

\bibitem{Tails} R. H. Price, Phys. Rev. D {\bf 5}, 2419 (1972);
C. Gundlach, R. H. Price, and J. Pullin, Phys. Rev. D {\bf 49}, 883
(1994); J. Bic\'ak, Gen. Relativ. Gravitation {\bf 3}, 331 (1972);
E. S. C. Ching, P. T. Leung, W. M. Suen, and K. Young, Phys. Rev.
Lett. {\bf 74}, 2414 (1995); E. S. C. Ching, P. T. Leung, W. M.
Suen, and K. Young, Phys. Rev. D {\bf 52}, 2118 (1995); S. Hod and
T. Piran, Phys. Rev. D {\bf 58}, 044018 (1998)
[arXiv:gr-qc/9801059]; S. Hod, Phys. Rev. D {\bf 58}, 104022 (1998)
[arXiv:gr-qc/9811032]; S. Hod, Phys. Rev. D {\bf 61}, 024033 (2000)
[arXiv:gr-qc/9902072]; S. Hod, Phys. Rev. D {\bf 61}, 064018 (2000)
[arXiv:gr-qc/9902073]; L. Barack, Phys. Rev. D {\bf 61}, 024026
(2000); S. Hod, Phys. Rev. Lett. {\bf 84}, 10 (2000)
[arXiv:gr-qc/9907096]; S. Hod, Phys. Rev. D {\bf 60}, 104053 (1999)
[arXiv:gr-qc/9907044]; S. Hod, Class. Quant. Grav. {\bf 26}, 028001
(2009) [arXiv:0902.0237]; S. Hod, Class. Quant. Grav. {\bf 18}, 1311
(2001) [arXiv:gr-qc/0008001]; S. Hod, Phys. Rev. D {\bf 66}, 024001
(2002) [arXiv:gr-qc/0201017]; R. J. Gleiser, R. H. Price, and J.
Pullin, Class. Quantum Grav. {\bf 25}, 072001 (2008); M. Tiglio, L.
E. Kidder, and S. A. Teukolsky, Class. Quant. Grav. {\bf 25}, 105022
(2008); R. Moderski and M. Rogatko, Phys. Rev. D {\bf 77}, 124007
(2008); X. He and J. Jing, Nucl. Phys.B {\bf 755}, 313 (2006); H.
Koyama and A. Tomimatsu, Phys. Rev. D {\bf 65}, 084031 (2002); B.
Wang, C. Molina, and E. Abdalla, Phys. Rev. D {\bf 63}, 084001
(2001); A. Zenginoglu and M. Tiglio, Phys. Rev.D {\bf 80}, 024044
(2009); A. J. Amsel, G.  T. Horowitz, D. Marolf, and M. M. Roberts,
J. High Energy Phys. 0909:044 (2009); S. Hod, Class. and Quant.
Grav. {\bf 30}, 237002 (2013) [arXiv:1402.4819]; S. Hod, Jour. of
High Energy Phys. {\bf 1309} (2013) 056 [arXiv:1310.4247].

\bibitem{Notenhc} It is important to stress the fact that existing
no-hair theorems (see \cite{Bek1} and references therein) rule out
the existence of asymptotically flat static charged
black-hole-scalar-field hairy configurations. In addition, it was
proved in \cite{Hodth} that stationary (or exponentially growing in
time) charged scalar fields (linearized bound-state charged scalar
configurations) cannot be supported in the external spacetime
regions of asymptotically flat charged Reissner-Nordstr\"om black
holes. Thus, charged scalar fields which propagate outside the outer
horizon of a spherically symmetric asymptotically flat
Reissner-Nordstr\"om black hole are expected to be swallowed by the
black hole or to be scattered away to infinity.

\bibitem{Hodth} S. Hod, Phys. Lett. B {\bf 713}, 505 (2012);
S. Hod, Phys. Lett. B {\bf 718}, 1489 (2013) [arXiv:1304.6474]; S.
Hod, Phys. Rev. D {\bf 91}, 044047 (2015) [arXiv:1504.00009].

\bibitem{Noteunit} We shall use natural units in which $G=c=\hbar=k_{\text{B}}=1$.

\bibitem{Notekj} It is important to emphasize that long-lived quasinormal resonances
of neutral massive scalar fields were observed in former numerical
studies \cite{Pjk} of composed black-hole-field systems. To the best
of our knowledge, our analysis (to be presented below) provides the
first fully {\it analytical} study of these intriguing long-lived
black-hole-field quasinormal resonances.

\bibitem{Pjk} A. Ohashi and M. Sakagami, Class. Quant.
Grav. {\bf 21}, 3973 (2004); R. A. Konoplya, A. V. Zhidenko, Phys.
Lett. B {\bf 609}, 377 (2005).

\bibitem{Noteqm} Note that $\mu$ and $q$ stand for $\mu/\hbar$ and $q/\hbar$, respectively.
Hence, these field parameters have the dimensions of
$($length$)^{-1}$.

\bibitem{Notegen} We shall assume $Q>0$ and $q>0$ without loss of
generality.

\bibitem{Chan} S. Chandrasekhar, {\it The Mathematical Theory of Black Holes}, (Oxford
University Press, New York, 1983).

\bibitem{HodPirpam} S. Hod and T. Piran, Phys. Rev. D {\bf 58},
024017 (1998) [arXiv:gr-qc/9712041]; S. Hod and T. Piran, Phys. Rev.
D {\bf 58}, 024018 (1998) [arXiv:gr-qc/9801001]; S. Hod and T.
Piran, Phys. Rev. D {\bf 58}, 024019 (1998) [arXiv:gr-qc/9801060].

\bibitem{Stro} T. Hartman, W. Song, and A. Strominger, JHEP 1003:118 (2010).

\bibitem{HodCQG2} S. Hod, Class. Quant. Grav. {\bf 23}, L23 (2006)
[arXiv:gr-qc/0511047].

\bibitem{Hodch1} S. Hod, Phys. Lett. A {\bf 374}, 2901 (2010)
[arXiv:1006.4439].

\bibitem{Hodch2} S. Hod, Phys. Lett. B {\bf 710}, 349 (2012)
[arXiv:1205.5087]; S. Hod, Phys. Lett. B {\bf 747}, 339 (2015)
[arXiv:1507.01943].

\bibitem{Konf} R. A. Konoplya, A. Zhidenko, Phys. Rev. D {\bf 88}, 024054
(2013).

\bibitem{Ric} M. Richartz and D. Giugno, Phys. Rev. D {\bf 90}, 124011 (2014).

\bibitem{Noteom} The integer parameters $m$ and $l\geq |m|$ are respectively the
azimuthal harmonic index and the spherical harmonic index which
characterize the scalar field mode. We shall henceforth omit these
harmonic indices for brevity.

\bibitem{Abram} M. Abramowitz and I. A. Stegun, {\it Handbook of
Mathematical Functions} (Dover Publications, New York, 1970).

\bibitem{Heun} A. Ronveaux, {\it Heun's differential equations} (Oxford University Press, Oxford, UK, 1995).

\bibitem{Notesy} It is important to note that the functions $S_{lm}(\theta)$
are the familiar associated Legendre polynomials
$Y_l^m(\cos\theta)$.

\bibitem{Detw} S. L. Detweiler, in Sources of Gravitational Radiation,
edited by L. Smarr (Cambridge University Press, Cambridge, England,
1979).

\bibitem{WKB1} B. F. Schutz and C. M. Will, Astrophys. J. {\bf 291}, L33 (1985).

\bibitem{WKB2} S. Iyer and C. M. Will, Phys. Rev. D {\bf 35}, 3621 (1987).

\bibitem{WKB3} S. Iyer, Phys. Rev. D {\bf 35}, 3632 (1987).

\bibitem{Will} L. E. Simone and C. M. Will, Class. Quant. Grav. {\bf 9}, 963 (1992).

\bibitem{Notewri} Note that we use the standard convention
$\omega=\omega_{\text{R}}-i\omega_{\text{I}}$ for the characteristic
quasinormal frequencies of the composed black-hole-field system.

\bibitem{Notemql} The requirement ${\bar\mu}<{\bar Q}$ stems from
the characteristic requirement $r_0>r_+$ for the peak location of
the effective scattering potential [see Eq. (\ref{Eq24})].

\bibitem{Notebnd} It is interesting to note that the relation
(\ref{Eq31}) with $r_0>r_+$ yields the characteristic inequality
$\Im\omega_0<\pi T_{\text{BH}}$ for the fundamental (least damped)
quasinormal resonance of the composed
RN-black-hole-charged-massive-scalar-field system. This
characteristic inequality implies that the relaxation time
$\tau_{\text{relax}}=1/\Im\omega_0$ of the composed black-hole-field
system is bounded from below by the relation
$\tau_{\text{relax}}\times T_{\text{BH}}>1/\pi$, in agreement with
the conjectured time-times-temperature (TTT) inequality
\cite{Hodtpi}.

\bibitem{Hodtpi} S. Hod, Phys. Rev. D {\bf 75}, 064013 (2007) [arXiv:gr-qc/0611004];
S. Hod, Class. Quant. Grav. {\bf 24}, 4235 (2007) [arXiv:0705.2306];
S. Hod, Phys. Rev. D {\bf 78}, 084035 (2008) [arXiv:0811.3806]; A.
Pesci, Class. Quant. Grav. {\bf 24}, 6219 (2007); S. Hod, Phys.
Lett. B {\bf 666}, 483 (2008) [arXiv:0810.5419]; S. Hod, Phys. Rev.
D {\bf 80}, 064004 (2009) [arXiv:0909.0314]; S. Hod, Phys. Rev. D.
{\bf 84}, 044046 (2011) [arXiv:1109.4080]; S. Hod, Phys. Lett. B
{\bf 715}, 348 (2012) [arXiv:1207.5282]; S. Hod, The Euro. Phys.
Jour. C {\bf 75}, L272 (2015) [arXiv:1410.2252].

\bibitem{Noterib} Note that $\omega_{\text{R}}>qQ/r_+$ whereas $\omega_{\text{I}}<2\pi
T_{\text{BH}}\cdot(n+1/2)<(n+1/2)/4M$ [see Eqs. (\ref{Eq28}) and
(\ref{Eq30})]. Thus, our assumption
$\omega_{\text{R}}\gg\omega_{\text{I}}$ [see (\ref{Eq21})] is valid
in the large coupling (eikonal) regime $qQ\gg n+1$ [see Eq.
(\ref{Eq2})].

\bibitem{NoteSchw} It is worth emphasizing the fact that vacuum polarization effects (the quantum
discharge phenomenon of charged black holes caused by the
Schwinger-type pair production mechanism \cite{Schw}) restrict the
physical parameters (masses and electric charges) of the composed
RN-black-hole-charged-massive-scalar-field system to the quantum
regime $qQ\ll\mu^2r^2_+/\hbar$ \cite{Hodch2}. Thus, one finds
$\sqrt{1-{\bar\mu}^2}qQ/\mu^2 r^2_0<qQ/\mu^2 r^2_+\ll1$ for the last
term in Eq. (\ref{Eq35}).

\bibitem{Schw} F. Sauter, Z. Phys. {\bf 69}, 742 (1931); W. Heisenberg and H. Euler,
Z. Phys. {\bf 98}, 714 (1936); J. Schwinger, Phys. Rev. {\bf 82},
664 (1951); M. A. Markov and V. P. Frolov, Teor. Mat. Fiz. {\bf 3},
3 (1970); W. T. Zaumen, Nature {\bf 247}, 531 (1974); B. Carter,
Phys. Rev. Lett. {\bf 33}, 558 (1974); G. W. Gibbons, Comm. Math.
Phys. {\bf 44}, 245 (1975); L. Parker and J. Tiomno, Astrophys.
Journ. {\bf 178}, 809 (1972); T. Damour and R. Ruffini, Phys. Rev.
Lett. {\bf 35}, 463 (1975); S. Hod, Phys. Rev. D {\bf 59}, 024014
(1999) [arXiv:gr-qc/9906004]; S. Hod and T. Piran, Gen. Rel. Grav.
{\bf 32}, 2333 (2000) [arXiv:gr-qc/0011003]; S. Hod, Phys. Lett. B
{\bf 693}, 339 (2010) [arXiv:1009.3695]; B.B. Levchenko,
arXiv:1208.0478 .

\bibitem{Noteap} Note that, in the extremal $(1-{\bar Q}^2)/(1-{\bar\mu}^2)\to 0\ (T_{\text{BH}}\to0)$ limit,
one finds from Eq. (\ref{Eq37}) the leading-order behavior
$\omega_n=qQ/r_+-i2\pi T_{\text{BH}}\cdot(n+1/2)$ for the
quasinormal resonant frequencies which characterize the composed
near-extremal RN-black-hole-charged-massive-scalar-field system.
This expression agrees with the results presented in \cite{Hodch1}.

\bibitem{Notepv} Note that, in the small mass ${\bar\mu}/{\bar Q}\ll1$ limit,
one finds from Eq. (\ref{Eq37}) the leading-order behavior
$\omega_n=qQ/r_+-i2\pi T_{\text{BH}}\cdot(n+1/2)$ for the
quasinormal resonant frequencies which characterize the composed
RN-black-hole-charged-scalar-field system. This expression agrees
with the results presented in \cite{Hodch2,Konf,Ric}.

\bibitem{Sham} M. O. P. Sampaio, C. Herdeiro, and M. Wang, Phys. Rev. D {\bf 90}, 064004
(2014).

\bibitem{Notesham} It is important to note that, for the case of quasi-bound states, a similar
behavior (that is, $M\omega_{\text{I}}\to0$ as $M\mu\to qQ$) was
observed numerically in the interesting work of Sampaio et. al.
\cite{Sham}. It is worth emphasizing that the quasi-bound states
studied in \cite{Sham} are characterized by the relation
$\omega_{\text{R}}<\mu$ ($qQ<M\mu$), whereas the free oscillations
(quasinormal resonances) analyzed in the present paper are
characterized by the relation $\mu>\omega_{\text{R}}$ ($M\mu<qQ$).
In addition, it is worth noting that the interesting work of Sampaio
et. al. \cite{Sham} uses {\it numerical} techniques in order to
study the quasi-bound states, whereas in the present work we use
{\it analytical} techniques (the WKB approximation) in order to
study the free oscillations of the charged massive fields.

\end{thebibliography}

\documentclass[preprintnumbers,amsmath,amssymbm,prd]{revtex4}
\usepackage{epsfig}
\usepackage{graphicx}

\begin{document}
\title{Slowly decaying resonances of charged massive scalar fields in the Reissner-Nordstr\"om black-hole spacetime}
\author{Shahar Hod}
\affiliation{The Ruppin Academic Center, Emeq Hefer 40250, Israel}
\affiliation{ }
\affiliation{The Hadassah Institute, Jerusalem 91010, Israel}
\date{\today}

\begin{abstract}
\ \ \ We determine the characteristic timescales associated with the
linearized relaxation dynamics of the composed
Reissner-Nordstr\"om-black-hole-charged-massive-scalar-field system.
To that end, the quasinormal resonant frequencies
$\{\omega_n(\mu,q,M,Q)\}_{n=0}^{n=\infty}$ which characterize the
dynamics of a charged scalar field of mass $\mu$ and charge coupling
constant $q$ in the charged Reissner-Nordstr\"om black-hole
spacetime of mass $M$ and electric charge $Q$ are determined {\it
analytically} in the eikonal regime $1\ll M\mu<qQ$. Interestingly,
we find that, for a given value of the dimensionless black-hole
electric charge $Q/M$, the imaginary part of the resonant
oscillation frequency is a monotonically {\it decreasing} function
of the dimensionless ratio $\mu/q$. In particular, it is shown that
the quasinormal resonance spectrum is characterized by the
asymptotic behavior $\Im\omega\to0$ in the limiting case $M\mu\to
qQ$. This intriguing finding implies that the composed
Reissner-Nordstr\"om-black-hole-charged-massive-scalar-field system
is characterized by extremely long relaxation times
$\tau_{\text{relax}}\equiv 1/\Im\omega\to\infty$ in the $M\mu/qQ\to
1^-$ limit.
\end{abstract}
\bigskip
\maketitle

\section{Introduction}

Wheeler's `no-hair' conjecture \cite{Whee,Car} presents a simple
physical picture according to which asymptotically flat black holes
in Einstein's theory of gravitation cannot support static matter
fields outside their horizons. Interestingly, various no-hair
theorems \cite{Whee,Car,Bek1,Chas,BekVec,Hart,Nun,Hod11} provide
strong support for the general validity of Wheeler's famous
conjecture. It is therefore expected that fundamental matter fields
that propagate in a static black-hole spacetime would eventually be
absorbed by the black hole or be scattered away to infinity. For
rotating black-hole spacetimes, a third possibility also exists
\cite{Hodrc,HerR}: thanks to the intriguing physical mechanism of
superradiant scattering of bosonic fields in spinning black-hole
spacetimes, these black holes can support stationary (rather than
static) linearized bound-state massive scalar configurations in
their external regions \cite{Hodrc}. Moreover, as explicitly shown
in \cite{HerR}, rotating black holes can also support genuine
bosonic hair (that is, non-linear stationary bosonic field
configurations) in their external regions. These rotating hairy
black-hole-bosonic-field configurations \cite{Hodrc,HerR} provide
explicit counterexamples to the no-hair conjecture in asymptotically
flat non-static spacetimes.

It is important to stress the fact that the elegant no-hair theorems
\cite{Whee,Car,Bek1,Chas,BekVec,Hart,Nun,Hod11}, which are valid for
asymptotically flat static black-hole configurations, say nothing
about the {\it timescale} associated with the dynamical process of
black-hole hair shedding \cite{Notetim}. This characteristic
relaxation timescale, $\tau_{\text{relax}}$, will be the main focus
of the present study.

The dynamics of fundamental matter and radiation fields in
black-hole spacetimes are characterized by {\it damped} quasinormal
oscillations of the form $e^{-i\omega t}$
\cite{QNMs,Notetails,Tails}. These exponentially decaying
oscillations are characterized by complex quasinormal resonant
frequencies $\{\omega_n\}_{n=0}^{n=\infty}$ whose values depend on
the physical parameters (such as mass, charge, angular momentum, and
intrinsic spin) of the composed black-hole-field system. In accord
with Wheeler's no-hair conjecture \cite{Whee,Car}, these
characteristic damped oscillations reflect the gradual decay of the
fields in the external regions of the black-hole spacetimes. In
particular, the characteristic timescale associated with the
relaxation dynamics of an external field in a black-hole spacetime
is determined by the imaginary part of the fundamental (least
damped) quasinormal resonant frequency which characterizes the
composed black-hole-field system:
\begin{equation}\label{Eq1}
\tau_{\text{relax}}\equiv 1/\Im\omega_0\  .
\end{equation}

The main goal of the present paper is to determine the
characteristic relaxation timescales, $\tau_{\text{relax}}$,
associated with the relaxation dynamics of charged massive scalar
fields in the charged Reissner-Nordstr\"om (RN) black-hole
spacetime. To that end, we shall explore below the quasinormal
resonance spectrum which characterizes the linearized relaxation
dynamics \cite{Notenhc,Hodth} of the composed
RN-black-hole-charged-massive-scalar-field system. As we shall show
below, the characteristic quasinormal resonances of this composed
black-hole-field system can be studied {\it analytically} in the
eikonal regime \cite{Noteunit}
\begin{equation}\label{Eq2}
1\ll M\mu<qQ\  ,
\end{equation}
where $\{q,\mu\}$ are respectively the charge coupling constant and
proper mass of the field, and $\{M,Q\}$ are respectively the mass
and electric charge of the RN black hole. In particular, below we
shall reveal the interesting fact that this composed
RN-black-hole-charged-massive-scalar-field system is characterized
by extremely long dynamical relaxation times,
$\tau_{\text{relax}}\equiv 1/\Im\omega_0\to\infty$, in the limiting
case $M\mu/qQ\to 1^-$ \cite{Notekj,Pjk}.

\section{Description of the system}

We shall analyze the quasinormal resonance spectrum which
characterizes the linearized relaxation dynamics of a scalar field
$\Psi$ of mass $\mu$ and charge coupling constant $q$ \cite{Noteqm}
in the spacetime of a Reissner-Nordstr\"om black hole of mass $M$
and electric charged $Q$ \cite{Notegen}. The curved black-hole
spacetime is described by the line element \cite{Chan}
\begin{equation}\label{Eq3}
ds^2=-f(r)dt^2+{1\over{f(r)}}dr^2+r^2(d\theta^2+\sin^2\theta
d\phi^2)\ ,
\end{equation}
where
\begin{equation}\label{Eq4}
f(r)=1-{{2M}\over{r}}+{{Q^2}\over{r^2}}\  .
\end{equation}
The zeros of the radial function $f(r)$,
\begin{equation}\label{Eq5}
r_{\pm}=M\pm(M^2-Q^2)^{1/2}\  ,
\end{equation}
determine the horizon radii of the charged RN black hole.

The familiar Klein-Gordon wave equation
\cite{HodPirpam,Stro,HodCQG2,Hodch1,Hodch2,Konf,Ric}
\begin{equation}\label{Eq6}
[(\nabla^\nu-iqA^\nu)(\nabla_{\nu}-iqA_{\nu})-\mu^2]\Psi=0\
\end{equation}
determines the linearized dynamics of the charged massive scalar
field in the curved RN black-hole spacetime, where
$A_{\nu}=-\delta_{\nu}^{0}{Q/r}$ is the electromagnetic potential of
the charged black hole. Substituting the field decomposition
\cite{Noteom}
\begin{equation}\label{Eq7}
\Psi(t,r,\theta,\phi)=\int\sum_{lm}e^{im\phi}S_{lm}(\theta)R_{lm}(r;\omega)e^{-i\omega
t} d\omega\ ,
\end{equation}
into the Klein-Gordon wave equation (\ref{Eq6}), and using the
black-hole metric function (\ref{Eq4}), one obtains
\cite{HodPirpam,Stro,HodCQG2} two ordinary differential equations of
the confluent Heun type \cite{Heun,Abram} for the eigenfunctions
$R(r)$ and $S(\theta)$ which respectively describe the radial and
angular behaviors of the charged massive scalar field in the curved
black-hole spacetime.

The ordinary differential equation which determines the spatial
behavior of the radial eigenfunction $R(r)$ is given by
\cite{HodPirpam,Stro,HodCQG2}
\begin{equation}\label{Eq8}
\Delta{{d} \over{dr}}\Big(\Delta{{dR}\over{dr}}\Big)+UR=0\ ,
\end{equation}
where $\Delta=r^2f(r)$, and
\begin{equation}\label{Eq9}
U=(\omega r^2-qQr)^2 -\Delta(\mu^2r^2+K_l)\  .
\end{equation}
Here $K_l=l(l+1)$ (with $l\geq |m|$) is the characteristic
eigenvalue of the angular eigenfunction $S_{lm}(\theta)$
\cite{Heun,Abram,Notesy}.

Defining the ``tortoise" radial coordinate $y$ by the differential
relation
\begin{equation}\label{Eq10}
dy={{dr}\over{f(r)}}\  ,
\end{equation}
and using the new radial eigenfunction
\begin{equation}\label{Eq11}
\psi=rR\  ,
\end{equation}
one can transform the radial equation (\ref{Eq8}) into the more
familiar form
\begin{equation}\label{Eq12}
{{d^2\psi}\over{dy^2}}+V\psi=0\
\end{equation}
of a Schr\"odinger-like ordinary differential equation, where the
effective radial potential in (\ref{Eq12}) is given by
\begin{equation}\label{Eq13}
V=V(r;M,Q,\omega,q,\mu,l)=\Big(\omega-{{qQ}\over{r}}\Big)^2-f(r)H(r)\
\end{equation}
with
\begin{equation}\label{Eq14}
H(r;M,Q,\mu,l)=
\mu^2+{{l(l+1)}\over{r^2}}+{{2M}\over{r^3}}-{{2Q^2}\over{r^4}}\ .
\end{equation}

The quasinormal resonant frequencies
$\{\omega_n(M,Q,\mu,q,l)\}_{n=0}^{n=\infty}$, which characterize the
linearized relaxation dynamics of the charged scalar field in the
charged black-hole spacetime, are determined by imposing on the
Schr\"odinger-like wave equation (\ref{Eq12}) the physically
motivated boundary conditions of purely ingoing waves at the
black-hole horizon and purely outgoing waves at spatial infinity
\cite{Detw}. That is,
\begin{equation}\label{Eq15}
\psi \sim
\begin{cases}
e^{-i (\omega-qQ/r_+)y} & \text{ as\ \ \ } r\rightarrow r_+\ \
(y\rightarrow -\infty)\ ; \\ y^{-iqQ}e^{i\sqrt{\omega^2-\mu^2} y} &
\text{ as\ \ \ } r\rightarrow\infty\ \ (y\rightarrow \infty)\  .
\end{cases}
\end{equation}

In the next section we shall study {\it analytically} the
quasinormal resonance spectrum
$\{\omega_n(M,Q,\mu,q,l)\}_{n=0}^{n=\infty}$ which characterizes the
relaxation dynamics of the composed
Reissner-Nordstr\"om-black-hole-charged-massive-scalar-field system
in the eikonal large-mass regime (\ref{Eq2}).

\section{The quasinormal resonance spectrum of the composed
Reissner-Nordstr\"om-black-hole-charged-massive-scalar-field system}

In the present section we shall perform a WKB analysis in order to
determine the complex resonant frequencies which characterize the
composed
Reissner-Nordstr\"om-black-hole-charged-massive-scalar-field system
in the large-mass regime
\begin{equation}\label{Eq16}
l+1\ll M\mu\  .
\end{equation}
In the eikonal large-mass regime (\ref{Eq16}), the radial potential
(\ref{Eq13}), which characterizes the dynamics of the charged
massive scalar field in the charged RN black-hole spacetime, can be
approximated by
\begin{equation}\label{Eq17}
V(r)=\Big(\omega-{{qQ}\over{r}}\Big)^2-\Big(1-{{2M}\over{r}}+{{Q^2}\over{r^2}}\Big)\mu^2\cdot\{1+O[(M\mu)^{-2}]\}.
\end{equation}
This radial potential has the form of an effective potential barrier
whose maximum $r_0$ is located at
\begin{equation}\label{Eq18}
r_0={{Q^2(q^2-\mu^2)}\over{qQ\omega-M\mu^2}}\  .
\end{equation}

As we shall now show, the fundamental complex resonances associated
with the effective scattering potential (\ref{Eq13}) can be
determined analytically in the large-mass regime (\ref{Eq16}) using
standard WKB methods \cite{WKB1,WKB2,WKB3,Will}. In particular, as
shown in \cite{WKB1,WKB2}, the WKB resonance condition which
characterizes the complex scattering resonances (the quasinormal
frequencies) of the Schr\"odinger-like radial equation (\ref{Eq12})
in the eikonal large-frequency regime is given by
\begin{equation}\label{Eq19}
{{iV_0}\over{\sqrt{2V^{(2)}_0}}}=n+{1\over 2}\ ,
\end{equation}
where the various derivatives $V^{(k)}_0\equiv d^{k}V/dy^{k}$ (with
$k\geq0$) that appear in the WKB resonance equation (\ref{Eq19}) are
evaluated at the maximum point $y=y_0(r_0)$ [see Eq. (\ref{Eq18})]
which characterizes the effective scattering potential $V(y)$.

Substituting Eqs. (\ref{Eq17}) and (\ref{Eq18}) into the WKB
resonance equation (\ref{Eq19}) and using the differential relation
(\ref{Eq10}), one finds the characteristic resonance condition
\begin{equation}\label{Eq20}
{{\Big(\omega-{{qQ}\over{r_0}}\Big)^2-\Big(1-{{2M}\over{r_0}}+{{Q^2}\over{r^2_0}}\Big)\mu^2}\over
{2\Big(1-{{2M}\over{r_0}}+{{Q^2}\over{r^2_0}}\Big)
\sqrt{{{qQ\omega-M\mu^2}\over{r^3_0}}}}}= -i\Big(n+{1\over2}\Big)\
\end{equation}
for the quasinormal resonant frequencies of the composed
Reissner-Nordstr\"om-black-hole-charged-massive-scalar-field system.
As we shall now show, the rather complicated resonance equation
(\ref{Eq20}) can be solved {\it analytically} in the regime
\cite{Notewri}
\begin{equation}\label{Eq21}
\omega_{\text{R}}\gg \omega_{\text{I}}\  .
\end{equation}
This strong inequality, which characterizes the fundamental
quasinormal frequencies of the composed black-hole-field system in
the eikonal regime (\ref{Eq2}) [see Eqs. (\ref{Eq25}) and
(\ref{Eq29}) below], enables one to {\it decouple} the real and
imaginary parts of the WKB resonance equation (\ref{Eq20}). In
particular, one finds
\begin{equation}\label{Eq22}
\Big(\omega_{\text{R}}-{{qQ}\over{r_0}}\Big)^2-\Big(1-{{2M}\over{r_0}}+{{Q^2}\over{r^2_0}}\Big)\mu^2=0\
\end{equation}
for the real part of the resonance equation (\ref{Eq20}), and
\begin{equation}\label{Eq23}
\Big(\omega_{\text{R}}-{{qQ}\over{r_0}}\Big)\omega_{\text{I}}=\Big(1-{{2M}\over{r_0}}+{{Q^2}\over{r^2_0}}\Big)
\sqrt{{{qQ\omega_{\text{R}}-M\mu^2}\over{r^3_0}}}\cdot\Big(n+{1\over2}\Big)\
\end{equation}
for the imaginary part of the resonance equation (\ref{Eq20}).

Substituting Eq. (\ref{Eq18}) into Eq. (\ref{Eq22}), one finds
\begin{equation}\label{Eq24}
{{r_0}\over{M}}={{1-{\bar\mu}^2+\sqrt{(1-{\bar\mu}^2)(1-{\bar
Q}^2)}}\over{1-({\bar\mu}/{\bar Q})^2}}
\end{equation}
and
\begin{equation}\label{Eq25}
\omega_{\text{R}}=q{\bar Q}\cdot {{1+\Big({{\bar\mu}\over{\bar
Q}}\Big)^2\sqrt{{{1-{\bar
Q}^2}\over{1-{\bar\mu}^2}}}}\over{1+\sqrt{{{1-{\bar
Q}^2}\over{1-{\bar\mu}^2}}}}}\  ,
\end{equation}
where
\begin{equation}\label{Eq26}
{\bar\mu}\equiv {{\mu}\over{q}}\ \ \ \ ; \ \ \ \ {\bar Q}\equiv
{{Q}\over{M}}\  .
\end{equation}
Note that the relations (\ref{Eq24}) and (\ref{Eq25}) are valid in
the regime ${\bar\mu}<{\bar Q}\leq1$ \cite{Notemql}, which
corresponds to
\begin{equation}\label{Eq27}
{{\mu}\over{q}}<{{Q}\over{M}}\leq 1\  .
\end{equation}

Interestingly, one finds from (\ref{Eq25}) that, for a given value
of the dimensionless black-hole electric charge ${\bar Q}$, the real
part of the resonant oscillation frequency, $\omega_{\text{R}}$, is
a monotonically {\it increasing} function of the dimensionless ratio
${\bar\mu}$. In particular, one finds the limiting behaviors [see
Eq. (\ref{Eq25})]
\begin{equation}\label{Eq28}
\{\omega_{\text{R}}\to {(qQ/r_+)}^+\ \ \ \text{for}\ \ \
{\bar\mu}/{\bar Q}\to0\}\ \ \ \ \ \text{and}\ \ \ \ \
\{\omega_{\text{R}}\to {(qQ/M)}^-\ \ \ \text{for}\ \ \
{\bar\mu}/{\bar Q}\to1^-\}\  .
\end{equation}

Substituting Eqs. (\ref{Eq24}) and (\ref{Eq25}) into Eq.
(\ref{Eq23}), one finds
\begin{equation}\label{Eq29}
M\omega_{\text{I}}=\sqrt{1-{\bar Q}^2}\Big[{{1-({\bar\mu}/{\bar
Q})^2}\over{1-{\bar\mu}^2+\sqrt{(1-{\bar\mu}^2)(1-{\bar
Q}^2)}}}\Big]^2\cdot\Big(n+{{1}\over{2}}\Big)
\end{equation}
for the imaginary parts of the quasinormal resonances which
characterize the composed RN-black-hole-charged-massive-scalar-field
system in the regime (\ref{Eq2}). Interestingly, one finds from
(\ref{Eq29}) that, for a given value of the dimensionless black-hole
electric charge ${\bar Q}$, the imaginary part of the resonant
oscillation frequency, $\omega_{\text{I}}$, is a monotonically {\it
decreasing} function of the dimensionless ratio ${\bar\mu}$. In
particular, one finds the limiting behaviors [see Eq. (\ref{Eq29})]
\begin{equation}\label{Eq30}
\{M\omega_{\text{I}}\to {{\sqrt{1-{\bar Q}^2}}\over{(1+\sqrt{1-{\bar
Q}^2})^2}}\cdot (n+1/2)\ \ \ \text{for}\ \ \ {\bar\mu}/{\bar
Q}\to0\}\ \ \ \ \ \text{and}\ \ \ \ \ \{M\omega_{\text{I}}\to 0^+ \
\ \ \text{for}\ \ \ {\bar\mu}/{\bar Q}\to1^-\}\  .
\end{equation}

Note that the expression (\ref{Eq29}) for the imaginary parts of the
RN-black-hole-charged-massive-scalar-field resonances can be written
in the compact form [see Eqs. (\ref{Eq5}) and (\ref{Eq24})]
\begin{equation}\label{Eq31}
\omega_{\text{I}}=2\pi
T_{\text{BH}}\Big({{r_+}\over{r_0}}\Big)^2\cdot\Big(n+{{1}\over{2}}\Big)\
,
\end{equation}
where
\begin{equation}\label{Eq32}
T_{\text{BH}}={{r_+-r_-}\over{4\pi r^2_+}}\
\end{equation}
is the Bekenstein-Hawking temperature of the RN black hole
\cite{Notebnd,Hodtpi,Noterib}.

\section{The regime of validity of the WKB approximation}

It is important to emphasize that our WKB results (\ref{Eq25}) and
(\ref{Eq29}) for the real and imaginary parts of the quasinormal
resonant frequencies which characterize the relaxation dynamics of
the charged massive scalar fields in the charged RN black-hole
spacetime were derived under the assumption that higher-order
correction terms that appear in the large-frequency WKB
approximation can be neglected. In particular, as explicitly shown
in \cite{WKB1,WKB2,WKB3,Will}, an extension of the WKB approximation
to include higher-order derivatives of the effective scattering
potential yields the correction term
\begin{equation}\label{Eq33}
\Lambda(n)={{1}\over{\sqrt{2V^{(2)}_0}}}\Big[{{1+(2n+1)^2}\over{32}}\cdot{{V^{(4)}_0}\over{V^{(2)}_0}}-
{{28+60(2n+1)^2}\over{1152}}\cdot\Big({{V^{(3)}_0}\over{V^{(2)}_0}}\Big)^2\Big]\
\end{equation}
on the r.h.s of the resonance equation (\ref{Eq19}). Thus, the WKB
resonance condition (\ref{Eq19}) is valid provided
\begin{equation}\label{Eq34}
{{\Lambda(n)}\over{n+{1\over 2}}}\ll1\  .
\end{equation}

Substituting Eqs. (\ref{Eq10}), (\ref{Eq17}), (\ref{Eq24}), and
(\ref{Eq25}) into Eq. (\ref{Eq34}), one finds
\begin{equation}\label{Eq35}
\Lambda(n)={{1}\over{qQ\sqrt{1-{\bar\mu}^2}}}\times
\Big[{1\over8}\Big(1-{{6M}\over{r_0}}+{{6Q^2}\over{r^2_0}}\Big)
-{3\over8}\Big(1-{{10M}\over{3r_0}}+{{10Q^2}\over{3r^2_0}}+{{4(1-{\bar\mu}^2)Q^2}
\over{3{\bar\mu}^2r^2_0}}\Big)\cdot(2n+1)^2\Big]\  ,
\end{equation}
which implies that the WKB resonance condition that we have used,
Eq. (\ref{Eq19}), is valid in the large coupling (eikonal) regime
[see Eq. (\ref{Eq34})] \cite{NoteSchw,Schw}
\begin{equation}\label{Eq36}
qQ\sqrt{1-{\bar\mu}^2}\gg n+{1\over2}\  .
\end{equation}

\section{Summary and Discussion}

We have determined the characteristic timescales associated with the
relaxation dynamics of the composed
Reissner-Nordstr\"om-black-hole-charged-massive-scalar-field system.
To that end, the quasinormal resonance spectrum
$\{\omega_n(\mu,q,M,Q)\}_{n=0}^{n=\infty}$ which characterizes the
dynamics of a linearized charged scalar field of mass $\mu$ and
charge coupling constant $q$ in the charged RN black-hole spacetime
of mass $M$ and electric charge $Q$ was studied {\it analytically}
in the eikonal regime $1\ll M\mu<qQ$ [see Eq. (\ref{Eq2})]. In
particular, we have derived the analytical expression [see Eqs.
(\ref{Eq25}) and (\ref{Eq29})]
\begin{equation}\label{Eq37}
M\omega(M,Q,\mu,q;n)=qQ\cdot {{1+\Big({{\bar\mu}\over{\bar
Q}}\Big)^2\sqrt{{{1-{\bar
Q}^2}\over{1-{\bar\mu}^2}}}}\over{1+\sqrt{{{1-{\bar
Q}^2}\over{1-{\bar\mu}^2}}}}}-i \sqrt{1-{\bar
Q}^2}\Big[{{1-({\bar\mu}/{\bar
Q})^2}\over{1-{\bar\mu}^2+\sqrt{(1-{\bar\mu}^2)(1-{\bar
Q}^2)}}}\Big]^2\cdot\Big(n+{{1}\over{2}}\Big)\
\end{equation}
for the quasinormal resonant frequencies which characterize the
composed RN-black-hole-charged-massive-scalar-field system in the
eikonal regime (\ref{Eq2}) \cite{Noteap,Notepv}.

The characteristic timescale $\tau_{\text{relax}}\equiv
1/\Im\omega_0$ associated with the linearized relaxation dynamics of
the composed RN-black-hole-charged-massive-scalar-field system is
determined by its fundamental (least damped) quasinormal resonant
frequency. In particular, from (\ref{Eq37}) one finds the expression
\begin{equation}\label{Eq38}
{{\tau_{\text{relax}}}\over{M}}={{2}\over{\sqrt{1-{\bar
Q}^2}}}\Big[{{1-{\bar\mu}^2+\sqrt{(1-{\bar\mu}^2)(1-{\bar
Q}^2)}}\over{1-({\bar\mu}/{\bar Q})^2}}\Big]^2
\end{equation}
for the characteristic relaxation time of the composed
black-hole-field system. Interestingly, one finds from (\ref{Eq38})
that the composed RN-black-hole-charged-massive-scalar-field system
is characterized by extremely long relaxation times in the limiting
case ${\bar\mu}/{\bar Q}\to 1^-$:
\begin{equation}\label{Eq39}
\tau_{\text{relax}}\to\infty\ \ \ \ \text{for}\ \ \ \
{{M\mu}\over{qQ}}\to 1^-\  .
\end{equation}
Thus, although Reissner-Nordstr\"om black holes cannot support
static matter fields outside their horizons
\cite{Bek1,Notenhc,Hodth,Sham,Notesham}, we conclude that these
black-hole spacetimes may host extremely {\it long-lived}
(exponentially decaying with long relaxation times,
$\tau_{\text{relax}}\gg M$) charged massive scalar fields in their
external regions.

\bigskip
\noindent
{\bf ACKNOWLEDGMENTS}
\bigskip

This research is supported by the Carmel Science Foundation. I would
like to thank Yael Oren, Arbel M. Ongo, Ayelet B. Lata, and Alona B.
Tea for helpful discussions.


\end{document}